OpenStudy (anonymous):

write y=3x^2+18x+25 in vertex form. Please explain how you do this.

jimthompson5910 (jim_thompson5910):

y=3x^2+18x+25 y-25=3x^2+18x y-25=3(x^2+6x) y-25=3(x^2+6x+9-9) y-25=3((x+3)^2-9) y-25=3(x+3)^2-27 y=3(x+3)^2-27+25 y=3(x+3)^2-2 The equation is now in vertex form y = a(x-h)^2 + k where a = 3, h = -3 and k = -2 So the vertex is (-3, -2)

OpenStudy (anonymous):

Where did the 9's come from in y-25=3(x^2+6x+9-9)?

jimthompson5910 (jim_thompson5910):

I took half of 6 in 6x to get 3. Then I squared 3 to get 9. I then added and subtracted 9 inside the parenthesis

OpenStudy (anonymous):

I have a test on this stuff tomorrow so i want to fully understand how to do the problems.

jimthompson5910 (jim_thompson5910):

I'm doing all this so I can get x^2+6x+9 which factors to (x+3)^2

jimthompson5910 (jim_thompson5910):

Post another and you do the steps and I'll see if you did it correctly or not

OpenStudy (anonymous):

Okay thanks a lot.

jimthompson5910 (jim_thompson5910):

yw

OpenStudy (anonymous):

wouldn't the 9's cancel out?

jimthompson5910 (jim_thompson5910):

they do (if you're going backward). The idea is that we don't want to change the right hand side So adding 9-9 (which is really 0) keeps things in balance

OpenStudy (anonymous):

Okay is there a simpler way to do this? lol

jimthompson5910 (jim_thompson5910):

the reason why i picked on 9 is because (x+3)^2 = x^2 + 6x + 9

jimthompson5910 (jim_thompson5910):

we have x^2+6x, which is missing that +9 but we can't just add 9 because that will change the equation, so we have to subtract 9 as well

jimthompson5910 (jim_thompson5910):

Here's an alternative way

jimthompson5910 (jim_thompson5910):

vertex form is y = a(x-h)^2 + k where (h,k) is the vertex

OpenStudy (anonymous):

Im sorry for being such a pain.

jimthompson5910 (jim_thompson5910):

To find h, which is the x coordinate of the vertex, use the formula h = -b/(2a)

jimthompson5910 (jim_thompson5910):

In the case of y=3x^2+18x+25, a = 3, b = 18 and c =25 So... h = -b/(2a) h = -18/(2*3) h = -18/6 h = -3

jimthompson5910 (jim_thompson5910):

We now can say that y = a(x-h)^2 + k becomes y = 3(x-(-3))^2 + k y = 3(x+3)^2 + k after we plug in a = 3 and h = -3

jimthompson5910 (jim_thompson5910):

with me so far?

OpenStudy (anonymous):

yeah so far

jimthompson5910 (jim_thompson5910):

to find k, the last piece of info, we just plug the x coordinate of the vertex in for x So plug in x = -3 into the original equation to get y=3x^2+18x+25 y=3(-3)^2+18(-3)+25 y=3(9)+18(-3)+25 y=27-54+25 y = -2 So the y coordinate of the vertex is y = -2 This means k = -2 So we can go from y = 3(x+3)^2 + k to y = 3(x+3)^2 + (-2) y = 3(x+3)^2 - 2

jimthompson5910 (jim_thompson5910):

The basic steps are: Step 1) Find the x coordinate of the vertex using the formula x = -b/(2a) This will give you the value of 'h' ------------------------------------ Step 2) Then use this x coordinate of the vertex to find the y coordinate of the vertex to obtain k. ------------------------------------ Step 3) Plug the values of 'a', h and k into y = a(x-h)^2 + k to complete the square and get the original equation into vertex form

OpenStudy (anonymous):

okay that helped me so much more. Thanks soooo much your awesome!

jimthompson5910 (jim_thompson5910):

you're welcome